HETAV PATEL

Grade 11,

The last two digits of 7^81 pls answer in an easy way

The last two digits of 7^81 pls answer in an easy way

Grade:12th pass

1 Answers

Deepak Kumar Shringi
6 years ago

Last 2digits of powers of7

7^1=07

7^2=49

7^3==43





7^8=01

So the last two digits repeat in cycles 01

Hence781781has the last two digits same as that of7171.So answer is :07